# Lennart Berg and Torben Wiedemann

### v1

@@ Abstract
Using the well-known realisation of the root system $H_{3}$ as a folding of $D_{6}$, one can construct examples of $H_{3}$-graded groups from Chevalley groups of type $D_{6}$. Such Chevalley groups are defined over a commutative ring $\mathcal{R}$, and the root groups of the resulting $H_{3}$-grading are coordinatised by $\mathcal{R}\times\mathcal{R}$. We show that every $H_{3}$-graded group arises as the folding of a $D_{6}$-graded group, or in other words, that it is coordinatised by $\mathcal{R}\times\mathcal{R}$ for some commutative ring $\mathcal{R}$.
@@ 1 Introduction
Motivated by these observations, we say that a $\Phi$-grading of a group $G$ is a family of non-trivial subgroups $(U_{\alpha})_{\alpha\in\Phi}$ satisfying some (purely combinatorial) properties of the root groups in Chevalley groups. Here $\Phi$ is any finite root system which is not assumed to be crystallographic or reduced. Our definition of root graded groups was introduced in [Wie23] and encompasses several existing concepts in the literature as special cases. In particular, it generalises Tits’ notion of RGD-systems (see [Tit92, p. 258]) and Shi’s notion of groups graded by a root system (see [Shi93]). An overview of the historic background of root graded groups is given in the preface of [Wie23].
@@
The main result of [Wie23] says that if $\Phi$ is irreducible, crystallographic and of rank at least $3$, then every $\Phi$-graded group is coordinatised by an algebraic structure (whose nature depends on $\Phi$). This algebraic structure is a division structure if and only if the corresponding $\Phi$-graded group is an RGD-system.
@@
In this paper, we prove a similar result for the non-crystallographic root systems $H_{3}$ and $H_{4}$, thereby extending the coordinatisation results of [Wie23] to the class of all irreducible finite root systems of rank at least 3. Specifically, we show in 8.31 that any root graded group $G$ of type $H_{3}$ or $H_{4}$ is coordinatised by $\mathcal{R}\times\mathcal{R}$ where $\mathcal{R}$ is a commutative ring (which depends on $G$). This is in line with Tits’ classical result from [Tit77, 3.5] that there exist no thick buildings of types $H_{3}$ and $H_{4}$: The ring $\mathcal{R}\times\mathcal{R}$ cannot be a division ring, and hence there do not exist RGD-systems of types $H_{3}$ and $H_{4}$. Most of the time, we can restrict our attention to the root system $H_{3}$, the transfer to $H_{4}$ being immediate.
@@
Since $H_{3}$ is not crystallographic, there are no Chevalley groups of type $H_{3}$, so it is not evident a priori that there actually exist $H_{3}$-graded groups. However, using the well-known realisation of the root system $H_{3}$ as a folding of $D_{6}$, examples of $H_{3}$-graded groups can be constructed from Chevalley groups of type $D_{6}$ (or, in fact, from any $D_{6}$-graded group). We will give a detailed description of this construction, which generalises the construction of the twisted Chevalley groups (see, for example, [Car72, Chapter 13]). In a similar way, one obtains examples of $H_{4}$-gradings as foldings of $E_{8}$-gradings. It is a consequence of our main result that, in fact, all root graded groups of types $H_{3}$ and $H_{4}$ arise in this way (see 8.40).
@@
The proof of our main result relies on the two main tools from [Wie23]: Firstly, the parametrisation theorem, which yields canonical isomorphisms between the root groups of an arbitrary $H_{3}$-graded groups $G$ and hence provides us with a parametrisation of the root groups by a single abelian group $(\mathcal{S},+)$. Secondly, the blueprint technique, which is a computational tool that allows us to equip $(\mathcal{S},+)$ with the structure of a commutative ring decomposing as $\mathcal{S}=\mathcal{S}_{1}\times\mathcal{S}_{2}$ for isomorphic ideals $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$. The blueprint technique is inspired by the work of Ronan-Tits on the construction of buildings in [RT87].
@@
In [MW23, 9.16], it is shown that an analogue of our coordinatisation result for $H_{3}$-graded groups holds for $H_{2}$-graded groups which arise from a Tits pentagon that satisfies a certain thickness condition. By [MW19, Section 5], the $H_{2}$-graded groups arising from a Tits pentagon are precisely those which satisfy a certain stability condition. Our main result is true regardless of any stability assumptions, but in return, it uses in a crucial way that the root systems $H_{3}$ and $H_{4}$ are of rank higher than 2.
@@
This paper is organised as follows. In Sections 2 and 3, we assemble some basic facts about root systems and introduce the notion of root graded groups. In Section 4, we construct examples of $H_{3}$-graded groups via foldings and explicitly compute the commutator formulas in these groups. The groundwork for the coordinatisation of $H_{3}$-graded groups is laid in Section 5, where we investigate the commutator relations and Weyl elements in such groups. The main result of this section is 5.20, which provides formulas for the action of squares of Weyl elements on all root groups. These formulas allow us to apply the parametrisation theorem, which is proven in [Wie23] for arbitrary root systems. In Section 6, we will formulate the parametrisation theorem for the root system $H_{3}$ and, for the reader’s convenience, sketch a simplified version of its proof in this special case. In Section 7, we will describe the general idea of the blueprint technique and then apply it to the special case of $H_{3}$-graded groups, which results in our coordinatisation theorem (Theorem 8.31). As a corollary, we obtain that every $H_{3}$-grading is the folding of a $D_{6}$-grading (Theorem 8.40).
@@ Acknowledgements
We are grateful to Bernhard Mühlherr for numerous insightful discussions and for his encouragement. The work of the first author was supported by the Studienstiftung des deutschen Volkes and the work of the second author was supported by the DFG Grant MU1281/7-1.
@@ 2 Preliminaries
We only consider unital rings in this paper, and homomorphism of rings preserve the identity. Further, all rings are assumed to be associative except when we specifically refer to “nonassociative rings”.
Throughout this paper, we will mostly use the standard terminology for finite root systems that are not necessarily crystallographic. However, though $H_{3}$ is a reduced root system, it will be practical to consider the non-reduced root system $GH_{3}$ (which we will introduce in 4.5) at some points. Hence we use the following definition of root systems.
@@
A root system is a finite subset $\Phi$ of a Euclidean space $(V,\cdot)$ such that $\Phi^{\sigma(\alpha)}=\Phi$ for all $\alpha\in\Phi$ where $\sigma(\alpha)\coloneqq\sigma_{\alpha}\colon V\rightarrow V,v\mapsto v-2\frac{\alpha\cdot v}{\alpha\cdot\alpha}\alpha$ denotes the reflection along $\alpha^{\perp}$. It is called reduced if $\Phi\cap\mathbb{R}\alpha=\{\alpha,-\alpha\}$ for all $\alpha\in\Phi$ and its Weyl group is $\operatorname{Weyl}(\Phi)\coloneqq\langle\sigma_{\alpha}\mid\alpha\in\Phi\rangle$.
Notation 2.5.
Let $\Phi$ be a (fixed) root system and let $A\subseteq\Phi$. Then $\Psi_{A}\coloneqq\langle A\rangle_{\mathbb{R}}\cap\Phi$ is called the root subsystem (of $\Phi$) spanned by $A$. In general, a root subsystem of $\Phi$ is the root subsystem spanned by some subset of $\Phi$. If $A$ has exactly two elements $\alpha$, $\beta$, we will write $\Psi_{\alpha,\beta}$ in place of $\Psi_{S}$. If $X_{n}$ is the type of a root system, we will often write $X_{n}$-subsystem” in place of “root subsystem of type $X_{n}$.
@@
Let $\Phi$ be a root system and let $A,B\subseteq\Phi$. We denote by $\mathord{]A,B[}$ (respectively, $\mathord{]A,B[}_{\text{cry}}$) the set of all roots $\gamma\in\Phi$ which can be written as $\gamma=\sum_{i=1}^{n}\lambda_{i}\alpha_{i}+\sum_{j=1}^{m}\mu_{j}\beta_{j}$ for $n,m\in\mathbb{Z}_{>0}$, $\alpha_{1},\ldots,\alpha_{n}\in A$, $\beta_{1},\ldots,\beta_{m}\in B$ and $\lambda_{1},\ldots,\lambda_{n},\mu_{1},\ldots,\mu_{m}$ positive real (respectively, integral) numbers. For non-proportional roots $\alpha,\beta\in\Phi$, the root interval between $\alpha$ and $\beta$ and the crystallographic root interval between $\alpha$ and $\beta$ are $\mathord{]\alpha,\beta[}\coloneqq\mathord{]\{\alpha\},\{\beta\}[}$ and $\mathord{]\alpha,\beta[}_{\text{cry}}\coloneqq\mathord{]\{\alpha\},\{\beta\}[}_{\text{cry}}$, and $\alpha,\beta$ are called (crystallographically) adjacent if $\mathord{]\alpha,\beta[}$ (respectively, $\mathord{]\alpha,\beta[}_{\text{cry}}$) is empty.
@@
In the literature, the root system $H_{2}$ is sometimes referred to as $I_{2}(5)$. Note that it is not crystallographic because $2\frac{\rho_{2}\cdot\rho_{3}}{\rho_{3}\cdot\rho_{3}}=2\cos(\frac{\pi}{5})=\frac{1+\sqrt{5}}{2}$ where $(\rho_{2},\rho_{3})$ denotes a root base of $H_{2}$. This number is known as the golden ratio, and we will denote it by $\tau$. Similarly, $H_{3}$ and $H_{4}$ are not crystallographic.
@@
Throughout this paper, we will make several assertions which follow from a straightforward but lengthy computation. The interested reader may find GAP [GAP21] code which verifies these assertions in [BW24].
@@
It follows from a straightforward inspection of $H_{3}$ that the assertion holds for $\alpha=\rho_{1}$ (see 2.11 for details). Then the general assertion follows from 2.9. ∎
@@
Let $\Phi$, $\Phi^{\prime}$ be root systems such that $\Phi^{\prime}$ is of rank 2. A pair $(\alpha,\beta)\in\Phi^{2}$ is called a $\Phi^{\prime}$-pair (in $\Phi$) if the root subsystem $\Psi_{\alpha,\beta}$ of $\Phi$ is of type $\Phi^{\prime}$ and $(\alpha,\beta)$ is a root base of this subsystem. If $(\alpha,\varepsilon)$ is an $H_{2}$-pair in $\Phi$, then the $H_{2}$-quintuple associated to $(\alpha,\varepsilon)$ is the tuple $(\alpha,\beta,\gamma,\delta,\varepsilon)$ where $\beta\coloneqq\tau\alpha+\varepsilon$, $\gamma\coloneqq\tau\alpha+\tau\varepsilon$ and $\delta\coloneqq\alpha+\tau\varepsilon$. Note that $(\beta,\gamma,\delta)$ is the interval ordering of $\mathord{]\alpha,\epsilon[}$ starting from $\alpha$. Similarly, the $A_{2}$-triple associated to an $A_{2}$-pair $(\alpha,\gamma)$ is $(\alpha,\alpha+\gamma,\gamma)$.

### v2

# Root graded groups of type $H_{3}$ and $H_{4}$

@@ Abstract
Using the well-known realisation of the root system $H_{4}$ as a folding of $E_{8}$, one can construct examples of $H_{4}$-graded groups from Chevalley groups of type $E_{8}$. Such Chevalley groups are defined over a commutative ring $\mathcal{R}$, and the root groups of the resulting $H_{4}$-grading are coordinatised by $\mathcal{R}\times\mathcal{R}$. We show that every $H_{4}$-graded group arises as the folding of an $E_{8}$-graded group, or in other words, that it is coordinatised by $\mathcal{R}\times\mathcal{R}$ for some commutative ring $\mathcal{R}$. We also prove similar assertions for $(D_{6},H_{3})$ in place of $(E_{8},H_{4})$.
@@ 1 Introduction
Motivated by these observations, we say that a $\Phi$-grading of a group $G$ is a family of non-trivial subgroups $(U_{\alpha})_{\alpha\in\Phi}$ satisfying some (purely combinatorial) properties of the root groups in Chevalley groups. Here $\Phi$ is any finite root system which is not assumed to be crystallographic or reduced. Our definition of root graded groups was introduced in [Wie24] and encompasses several existing concepts in the literature as special cases. In particular, it generalises Tits’ notion of RGD-systems (see [Tit92, p. 258]) and Shi’s notion of groups graded by a root system (see [Shi93]). An overview of the historic background of root graded groups is given in the preface of [Wie24].
@@
The main result of [Wie24] says that if $\Phi$ is irreducible, crystallographic and of rank at least $3$, then every $\Phi$-graded group is coordinatised by an algebraic structure (whose nature depends on $\Phi$). This algebraic structure is a division structure if and only if the corresponding $\Phi$-graded group is an RGD-system.
@@
In this paper, we prove a similar result for the non-crystallographic root systems $H_{3}$ and $H_{4}$, thereby extending the coordinatisation results of [Wie24] to the class of all irreducible finite root systems of rank at least 3. Specifically, we show in 9.33 that any root graded group $G$ of type $H_{3}$ or $H_{4}$ is coordinatised by $\mathcal{R}\times\mathcal{R}$ where $\mathcal{R}$ is a commutative ring (which depends on $G$). Since $\mathcal{R}\times\mathcal{R}$ cannot be a division ring, it follows that there do not exist RGD-systems of types $H_{3}$ or $H_{4}$. This is a classical result of Tits [Tit77, Hauptsatz], who proved it by showing that there are no RGD-systems of type $H_{2}$ [Tit76, Theorème 1].
@@
In the following, let $\ell\in\{3,4\}$. Since $H_{\ell}$ is not crystallographic, there are no Chevalley groups of type $H_{\ell}$, so it is not evident a priori that there actually exist $H_{\ell}$-graded groups. However, using the well-known realisation of the root system $H_{4}$ as a folding of $E_{8}$, examples of $H_{4}$-graded groups can be constructed from $E_{8}$-graded groups (such as, for example, Chevalley groups of type $E_{8}$). We will give a detailed description of this construction, which generalises the construction of the twisted Chevalley groups (see, for example, [Car72, Chapter 13]). In a similar way, one obtains examples of $H_{3}$-gradings as foldings of $D_{6}$-gradings. It is a consequence of our main result that, in fact, all root gradings of type $H_{3}$ or $H_{4}$ arise as foldings of $D_{6}$- or $E_{8}$-gradings (see 10.11).
@@
The proof of our main result relies on the two main tools from [Wie24]: The first one is the parametrisation theorem, which yields canonical isomorphisms between the root groups of an arbitrary $H_{\ell}$-graded group $G$ and hence provides us with a parametrisation of the root groups by a single abelian group $(\mathcal{S},+)$. Our second main tool is the blueprint technique, which is a computational method that allows us to equip $(\mathcal{S},+)$ with the structure of a commutative ring decomposing as $\mathcal{S}=\mathcal{S}_{1}\times\mathcal{S}_{2}$ for isomorphic ideals $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$. The blueprint technique follows the ideas of Ronan-Tits on the construction of buildings in [RT87], but in a “reversed” setting.
@@
In [MW24, 9.16], it is shown that an analogue of our coordinatisation result for $H_{\ell}$-graded groups holds for $H_{2}$-graded groups which arise from a Tits pentagon that satisfies a certain thickness condition. By [MW19, Section 5], the $H_{2}$-graded groups arising from a Tits pentagon are precisely those which satisfy a certain stability condition. Our main result is true regardless of any stability assumptions, but in return, it uses in a crucial way that the root systems $H_{3}$ and $H_{4}$ are of rank higher than 2.
@@
This paper is organised as follows. In Sections 2 and 3, we assemble some basic facts about root systems and introduce the notion of root graded groups. In Section 4, we construct examples of $H_{\ell}$-graded groups via foldings and explicitly compute the commutator formulas in these groups. The groundwork for the coordinatisation of $H_{\ell}$-graded groups is laid in Section 5, where we investigate the commutator relations and Weyl elements in such groups. The main result of this section is 5.20, which provides formulas for the action of squares of Weyl elements on all root groups. These formulas allow us to apply the parametrisation theorem, which is proven in [Wie24] for arbitrary root systems. In Section 6, we formulate the parametrisation theorem for the root systems $H_{3}$ and $H_{4}$ and, for the reader’s convenience, sketch a simplified version of its proof in this special case. We also define the notion of coordinatisations of $H_{\ell}$-graded groups, which is studied in more detail in Section 7. In Section 8, we describe the general idea of the blueprint technique and then, in Section 9, apply it to the special case of $H_{\ell}$-graded groups. This leads to our main result, the coordinatisation theorem (Theorem 9.33). As a corollary, we obtain in Section 10 that every root grading of type $H_{3}$ or $H_{4}$ is the folding of a grading of type $D_{6}$ or $E_{8}$, respectively (Theorem 10.11).
@@ Acknowledgements
We are grateful to Bernhard Mühlherr for numerous insightful discussions and for his encouragement. We also want to thank the referee for carefully reading our paper and for making many valuable suggestions. The work of the first author was supported by the Studienstiftung des deutschen Volkes and the work of the second author was supported by the DFG Grant MU1281/7-1.
@@ 2 Preliminaries
We only consider rings with identity in this paper, and homomorphism of rings preserve the identity. Further, all rings are assumed to be associative except when we specifically refer to “nonassociative rings”.
Throughout this paper, we will mostly use the standard terminology for finite root systems that are not necessarily crystallographic. However, though $H_{3}$ and $H_{4}$ are reduced root systems, it will be practical to consider the non-reduced root systems $GH_{3}$ and $GH_{4}$ (which we will introduce in 2.16) at some points. Hence we use the following definition of root systems.
@@
A root system is a finite subset $\Phi$ of a Euclidean space $(V,\cdot)$ such that $\Phi^{\sigma(\alpha)}=\Phi$ for all $\alpha\in\Phi$ where
$\sigma(\alpha)\coloneqq\sigma_{\alpha}\colon V\rightarrow V,v\mapsto v^{\sigma(\alpha)}\coloneqq v-2\frac{\alpha\cdot v}{\alpha\cdot\alpha}\alpha$
denotes the reflection along $\alpha^{\perp}$. It is called reduced if $\Phi\cap\mathbb{R}\alpha=\{\alpha,-\alpha\}$ for all $\alpha\in\Phi$ and its Weyl group is $\operatorname{Weyl}(\Phi)\coloneqq\langle\sigma_{\alpha}\mid\alpha\in\Phi\rangle$. It is called crystallographic if for all $\alpha,\beta\in\Phi$, the number $2\frac{\alpha\cdot\beta}{\beta\cdot\beta}$ is an integer.
Remark 2.5.
Note our convention in 2.4 that reflections act on $V$ from the right. Consequently, we have $\sigma_{v^{w}}=\sigma_{v}^{w}$ for all $v\in V\setminus\{0\}$ and all orthogonal $w\in\operatorname{GL}(V)$, and not the formula $\sigma_{w(v)}=w\circ\sigma_{v}\circ w^{-1}$ that holds for left actions.
Notation 2.6.
Let $\Phi$ be a (fixed) root system and let $A\subseteq\Phi$. Then $\Psi_{A}\coloneqq\langle A\rangle_{\mathbb{R}}\cap\Phi$ is called the root subsystem (of $\Phi$) spanned by $A$. In general, a root subsystem of $\Phi$ is the root subsystem spanned by some subset of $\Phi$. If $A=\{\alpha_{1},\ldots,\alpha_{n}\}$, we will write $\Psi_{\alpha_{1},\ldots,\alpha_{n}}$ for $\Psi_{A}$. If $X_{n}$ is the type of a root system, we will often write $X_{n}$-subsystem” in place of “root subsystem of type $X_{n}$.
Definition 2.7.
Let $\Phi$, $\Phi^{\prime}$ be root systems such that $\Phi^{\prime}$ is of rank 2. A pair $(\alpha,\beta)\in\Phi^{2}$ is called a $\Phi^{\prime}$-pair (in $\Phi$) if the root subsystem $\Psi_{\alpha,\beta}$ of $\Phi$ is of type $\Phi^{\prime}$ and $(\alpha,\beta)$ is a root base of this subsystem.
Notation 2.8.
Let $\Phi$ be a root system, let $\Delta$ be a root base of $\Phi$ and denote by $\Pi$ the corresponding positive subsystem of $\Phi$. We put
$E_{2}^{+}(\Delta)\coloneqq E_{2}^{+}(\Delta,\Phi)\coloneqq\Phi\cap\{a\delta+b\delta^{\prime}\mid a,b\in\mathbb{R}_{\geq 0},\delta,\delta^{\prime}\in\Delta\}=\Pi\cap\bigcup_{\delta,\delta^{\prime}\in\Delta}\Psi_{\delta,\delta^{\prime}}.$
Remark 2.9.
Let $\Phi$ be a root system, let $\Delta$ be a root base of $\Phi$ and let $\alpha,\beta\in\Phi$ be non-proportional. Then there exists $w\in\operatorname{Weyl}(\Phi)$ such that $\alpha^{w}$ and $\beta^{w}$ both lie in $E_{2}^{+}(\Delta,\Phi)$. For this reason, we will only describe the commutator relations in $H_{\ell}$-graded groups (for $\ell\in\{3,4\}$) for pairs of roots in $E_{2}^{+}(\Delta,H_{\ell})$ where $\Delta$ is some fixed root base of $H_{\ell}$.
Notation 2.10.
Let $\Phi$ be a root system of rank $n$ and let $\Delta=(\delta_{1},\ldots,\delta_{n})$ be an ordered root base of $\Phi$. For all $a\in\mathbb{R}^{n}$, we put $[a_{1},\ldots,a_{n}]_{\Delta}\coloneqq\sum_{i=1}^{n}a_{i}\delta_{i}$. When $\Delta$ is clear from the context, we will simply write $[a_{1},\ldots,a_{n}]$ for $[a_{1},\ldots,a_{n}]_{\Delta}$.
@@
Let $\Phi$ be a root system and let $\alpha,\beta\in\Phi$ be non-proportional. The root interval between $\alpha$ and $\beta$ and the crystallographic root interval between $\alpha$ and $\beta$ are
$\mathord{]\alpha,\beta[}\coloneqq\Phi\cap\{a\alpha+b\beta\mid a,b\in\mathbb{R}_{>0}\},\quad\mathord{]\alpha,\beta[}_{\text{cry}}\coloneqq\Phi\cap\{a\alpha+b\beta\mid a,b\in\mathbb{Z}_{>0}\},$
respectively. Further, $\alpha,\beta$ are called (crystallographically) adjacent if $\mathord{]\alpha,\beta[}$ (respectively, $\mathord{]\alpha,\beta[}_{\text{cry}}$) is empty.
@@
The cardinalities of the root systems $H_{2}$, $H_{3}$ and $H_{4}$ are $10$, $30$ and $120$, respectively. Their Weyl groups have respective orders $10$, $120$ and $120\cdot 120$. Further, a brief computation shows that $\operatorname{Weyl}(H_{\ell})$ acts transitively on $H_{\ell}$ for $\ell=2$, which implies that the same holds for $\ell\in\{3,4\}$ as well.
@@
In the literature, the root system $H_{2}$ is sometimes referred to as $I_{2}(5)$. Note that it is not crystallographic: If $(\rho_{2},\rho_{3})$ denotes a root base of $H_{2}$, then $2\frac{\rho_{2}\cdot\rho_{3}}{\rho_{3}\cdot\rho_{3}}=2\cos(\frac{\pi}{5})=\frac{1+\sqrt{5}}{2}$ is not an integer. This number is known as the golden ratio, and we will denote it by $\tau$ throughout the paper. Similarly, $H_{3}$ and $H_{4}$ are not crystallographic because they contain $H_{2}$.
Notation 2.16.
Let $\ell\in\{2,3,4\}$. The root system $GH_{\ell}$ is $GH_{\ell}\coloneqq H_{\ell}\cup\tau H_{\ell}$ where $\tau\coloneqq\frac{1+\sqrt{5}}{2}$ denotes the golden ratio. We also call it the golden version of $H_{\ell}$. It is a non-reduced root system with the same Weyl group as $H_{\ell}$ and with two roots in each ray, whose lengths differ by a factor of $\tau$. When we speak of $H_{\ell}$ as a subset of $GH_{\ell}$, we will always mean the set of indivisible roots in $GH_{\ell}$ (which is not a subsystem in the sense of 2.6). A root base of $GH_{\ell}$ (in standard order) is a root base of the subset $H_{\ell}$ (in standard order). The map $\phi\colon GH_{\ell}\rightarrow H_{\ell},\alpha\mapsto\frac{1}{\lVert\alpha\rVert}\alpha$ which sends each root in $GH_{\ell}$ to the indivisible root in its ray is called the scaling map.
Remark 2.17.
See Figure 1 for a graphic depiction of $GH_{2}$ and [MW24] for more details on $GH_{2}$, which inspired our definition of $GH_{3}$ and $GH_{4}$.
Remark 2.18.
Throughout this paper, we will make several assertions which follow from a straightforward but lengthy computation. The interested reader may find GAP [GAP24] code which verifies these assertions in [BW25].
@@
By 2.14, we can choose $\rho_{2},\rho_{3}\in H_{3}$ such that $(\alpha,\rho_{2},\rho_{3})$ is a root base of $H_{3}$ in standard order. Now the assertion follows from a straightforward inspection of $H_{3}$ (see 2.18 for details). ∎
@@
If $(\alpha,\varepsilon)$ is an $H_{2}$-pair in $\Phi$, then the $H_{2}$-quintuple associated to $(\alpha,\varepsilon)$ is the tuple $(\alpha,\beta,\gamma,\delta,\varepsilon)$ where $\beta\coloneqq\tau\alpha+\varepsilon$, $\gamma\coloneqq\tau\alpha+\tau\varepsilon$ and $\delta\coloneqq\alpha+\tau\varepsilon$. Note that $(\beta,\gamma,\delta)$ is the interval ordering of $\mathord{]\alpha,\varepsilon[}$ starting from $\alpha$. Similarly, the $A_{2}$-triple associated to an $A_{2}$-pair $(\alpha,\gamma)$ is $(\alpha,\alpha+\gamma,\gamma)$.